# MemEvo: Memory-Evolving Incremental Multi-view Clustering

### v1

@@ Abstract
Incremental multi-view clustering aims to achieve stable clustering results while addressing the stability-plasticity dilemma (SPD) in incremental views. At the core of SPD is the challenge that the model must have enough plasticity to quickly adapt to new data, while maintaining sufficient stability to consolidate long-term knowledge and prevent catastrophic forgetting. Inspired by the hippocampal-prefrontal cortex collaborative memory mechanism in neuroscience, we propose a Memory-Evolving Incremental Multi-view Clustering method (MemEvo) to achieve this balance. First, we propose a hippocampus-inspired view alignment module that captures the gain information of new views by aligning structures in continuous representations. Second, we introduce a cognitive forgetting mechanism that simulates the decay patterns of human memory to modulate the weights of historical knowledge. Additionally, we design a prefrontal cortex-inspired knowledge consolidation memory module that leverages temporal tensor stability to gradually consolidate historical knowledge. By integrating these modules, MemEvo achieves strong knowledge retention capabilities in scenarios with a growing number of views. Extensive experiments demonstrate that MemEvo exhibits remarkable advantages over existing state-of-the-art methods.
Introduction
@@
Recent advancements in multimedia technology have significantly transformed data analysis paradigms, where Multi-View Clustering (MVC) has demonstrated notable advantages in unsupervised scenarios by integrating heterogeneous data sources (Du et al. 2025; Wen et al. 2024; Fu et al. 2024). Traditional MVC methods have achieved significant success in processing static, complete datasets, as they require a complete set of views to achieve effective knowledge fusion (Du et al. 2024). However, data is not static but evolves dynamically over time in many real-world applications. For example, medical monitoring may extend from clinical measurements to imaging scan data, and multi-view data facilitates tracking of patient progress. This incremental property poses a major challenge to traditional methods, driving incremental multi-view clustering (IMVC) to become a key research direction. IMVC aims to efficiently update clustering results when new views arrive, thereby avoiding retraining the model from scratch. However, its core challenge is to resolve a fundamental contradiction: the Stability-Plasticity Dilemma (SPD) (Wu, Gong, and Li 2021). On the one hand, the model must possess sufficient plasticity to enable rapid learning and adaptation to new patterns and knowledge contained in new data. If the model is too rigid, it will be unable to capture dynamic changes in the data, leading to a rapid decline in performance. On the other hand, the model must also maintain sufficient stability to consolidate long-term accumulated knowledge and avoid catastrophic forgetting when learning new information. Therefore, an ideal IMVC model must achieve a balance between these two aspects.
@@
Recent incremental multi-view clustering (IMVC) methods have shown significant progress. For example, Wan et al. (Wan et al. 2022) dynamically update a consensus partition matrix by transferring knowledge from historical views, enabling incremental learning. Yan et al. (Yan et al. 2024) further introduced a category memory library to store learned categories, while Qu et al. (Qu et al. 2024) proposed a lightweight framework that retains and jointly optimizes the consensus anchor graph of previous views. Despite existing IMVC methods having made valuable explorations, existing IMVC methods often address only specific aspects of SPD. Some prioritize plasticity, rapidly integrating new data at the risk of catastrophic forgetting, while others emphasize stability through regularization, potentially hindering adaptation to new views. A key limitation is that existing methods fail to balance old knowledge retention and new knowledge integration, which causes performance bottlenecks as views continuously grow. To address this dilemma, we drew inspiration from the memory system of the human brain. Cognitive science research has found that the human brain, particularly the collaborative memory system between the hippocampus and the prefrontal cortex, serves as a paradigm for balancing plasticity and stability. The hippocampus can quickly encode new episodic memories and is highly plastic (Arani, Sarfraz, and Zonooz 2022). In contrast, the prefrontal cortex integrates and consolidates these temporary memories into long-term knowledge (highly stable) through a slow, gradual process. This complementary mechanism of rapid learning and slow consolidation enables humans to continue learning without easily forgetting the past. Furthermore, the acquisition of knowledge is not merely an accumulation of information. The brain also employs a complex active adaptive forgetting mechanism to discard outdated information and maintain cognitive efficiency (Anderson and Hulbert 2021). This forgetting is not a passive process of memory decay but an active process of neural remodeling that occurs immediately after learning behavior is completed (Hardt, Nader, and Nadel 2013).
@@
Inspired by this, we propose a Memory-Evolving Incremental Multi-view Clustering method (MemEvo) that aims to solve the plasticity-stability dilemma. We should clarify that MemEvo is not designed to create a biologically accurate brain simulation. Instead, we borrow from their division of roles between plastic learning and stable integration to build a more effective incremental clustering algorithm. To this end, we have designed a collaborative framework comprising three core modules. First, we propose a hippocampus-inspired view alignment module, which effectively captures incremental information from new views by aligning structures in the continuous representation space, thereby ensuring plasticity of our model. Second, we introduce a cognitive forgetting mechanism. This mechanism simulates the decay patterns of human memory, selectively forgetting historical information by adjusting weights. Finally, we designed a prefrontal cortex-inspired knowledge consolidation memory module, which leverages the stability of temporal tensors to gradually consolidate historical knowledge, ensuring the stability of the model. By combining these three core modules, MemEvo achieves remarkable knowledge retention capabilities in view-incremental scenes. Our main contributions are summarized as follows:
@@
  • A Novel Neuroscience-Inspired Framework: This paper first introduces the hippocampus-prefrontal cortex collaborative memory model into incremental multi-view clustering, providing a new perspective for solving the stability-plasticity dilemma.
  • Collaborative Module Design: We designed a collaborative optimization framework that organically integrates three core modules: view alignment, cognitive forgetting, and knowledge consolidation memory, achieving efficient and stable incremental learning.
@@
  • State-of-the-Art Performance: Extensive experiments show that our proposed method outperforms existing incremental clustering methods.
Related Work
Incremental multi-view clustering
@@
Incremental Multi-View Clustering (IMVC) continuously integrates new views while retaining historical knowledge, thereby achieving stable clustering. Existing literature achieves incremental updates by maintaining historical knowledge carriers (Wan et al. 2022; Lu et al. 2024; Wan et al. 2024; Chen et al. 2025). For example, CMVC (Wan et al. 2022) and CAC (Yan et al. 2024) maintain consensus partition matrices to achieve knowledge reuse. FDMVC (Lu et al. 2024) maintains semantically consistent representations for updates. LAIMVC (Qu et al. 2024) maintains consensus anchor graphs and aligns new view anchors through permutation matrices. Consistent with existing research, MemEvo is based on the following assumptions: each view contains complete class information, and class growth scenarios are not considered. Based on this, we focus on how to simulate cognitive mechanisms to achieve efficient capture of multi-view information.
Nonconvex Tensor rank substitution
Non-convex methods have been proven (Zhong et al. 2015) to approximate the tensor rank more accurately, and thus have become the current dominant technical paradigm in traditional multi-view clustering (Guo et al. 2022; Long et al. 2025; Ji and Feng 2025). Let $\{\mathbf{X}_{t}\}_{t=1}^{T}\in\mathbb{R}^{n\times d_{t}}$ be multi-view data with $T$ views, where $d_{t}$ is the dimension of the view and $n$ is the sample number. We can optimize the following objective function to achieve our goal.
$\min_{\mathbf{A}_{t},\mathbf{Z}_{t}}\sum_{t=1}^{T}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1}+\lambda\|\mathcal{Z}\|_{low},s.t.\mathcal{Z}=\Phi(\mathbf{Z}_{1},..,\mathbf{Z}_{t})$ (1)
@@
where $\mathbf{A}_{t}\in\mathbb{R}^{m\times d_{t}}$ (where $m$ is the latent space dimension) is the basic matrix and $\mathbf{Z}_{t}$ is the representation for $t$-view. Besides, $\Phi(\cdot)$ constructs the tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ by stacking representation $\mathbf{Z}_{t}\in\mathbb{R}^{n\times m}$ from each view into a three-order tensor, $\|\cdot\|_{low}$ is the rank constraint term. Here, a commonly used nonconvex substitution function (Lu et al. 2015) is introduced, which is defined as:
@@
$\|\mathcal{Z}\|_{ARMR}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\sum_{i=1}^{min(n_{1},n_{2})}\left(\frac{1-e^{-\mathcal{S}_{f}^{k}(i,i)}}{1+e^{-\mathcal{S}_{f}^{k}(i,i)}}\right)$ (2)
@@
With the Fast Fourier Transform (FFT) along the third dimension, we can obtain $\mathcal{Z}_{f}^{k}$ by $fft(\mathcal{Z})=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\mathcal{Z}_{f}^{k}$. Then, we perform tensor singular value decomposition(t-SVD) to obtain $\mathcal{Z}_{f}^{k}=\mathcal{U}*\mathcal{S}_{f}^{k}*\mathcal{V}^{T}$ (Kilmer et al. 2013).
@@
Figure 1: The framework of MemEvo. The model achieves the retention of historical knowledge and the learning of new knowledge through three core modules: cognitive forgetting, view alignment, and knowledge consolidation.
The Proposed method
Overview and Motivation
@@
In incremental multi-view clustering, maintaining a balance between learning new knowledge (plasticity) and retaining existing knowledge (stability) is a major challenge. To address this, we draw inspiration from the human brain’s memory system. Specifically, we have simulated the collaborative memory mechanism between the hippocampus and the prefrontal cortex. As shown in Figure 1, MemEvo contains three core components organized to address the Stability-Plasticity Dilemma: a hippocampus-inspired view alignment module, a cognitive forgetting mechanism, and a prefrontal cortex-inspired knowledge consolidation memory module.
Plasticity: New Knowledge Encoding & Alignment
@@
When new view data arrives, the model must quickly learn and adapt to the new knowledge it contains. We simulate the rapid learning ability of the brain through incremental representation learning and view alignment.
@@
Foundation: Incremental Representation Learning. First, we designed an incremental representation learning strategy to establish a foundation for subsequent memory operations. When a new data view $\mathbf{X}_{t}$ is received, we reconstruct the original data by a representation $\mathbf{Z}_{t}$ and an orthogonal basis matrix $\mathbf{A}_{t}$. This process can be expressed as reconstruction loss:
$\mathcal{L}_{recon}=\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1},\quad s.t.\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I}.$ (3)
@@
where $\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I}$ ensures that the basis vectors are uncorrelated (Chen et al. 2022). $\|\cdot\|_{2,1}$ is the $\ell_{2,1}$ norm, which is used to enhance the robustness of model (Nie et al. 2010).
@@
View Alignment Module (VAM): Simulating the rapid association of the hippocampus. In cognitive science, the hippocampus serves as a temporary storage center for short-term memory, playing a key role in the initial encoding and integration of new and old information (McClelland, McNaughton, and Lampinen 2020). To simulate this rapid initial association, we propose a hippocampus-inspired view alignment module (VAM). This module aims to achieve initial interaction between new and old knowledge, enabling the model to adapt to past knowledge while learning new tasks. Specifically, we introduce an orthogonal matrix $\mathbf{P}_{t}\in\mathbb{R}^{m\times m}$ to align the current representation $\mathbf{Z}_{t}$ with the previous representation $\mathbf{Z}_{t-1}$. This process can be expressed as
$\mathcal{L}_{align}=\|\mathbf{Z}_{t}-{\mathbf{Z}_{t-1}\mathbf{P}_{t}}\|_{F}^{2},s.t.\mathbf{P}_{t}\mathbf{P}_{t}^{T}=\mathbf{I}.$ (4)
Although VAM uses a technically mature orthogonal alignment method, this choice is a principled abstraction of hippocampal. The hippocampus is characterized by its ability to rapidly encode new events with high plasticity. Our VAM module aligns the current view with the previous view, enabling rapid processing, which corresponds to the encoding function of the hippocampus.
Stability: Knowledge Evolution and Consolidation
@@
Although plasticity enables systems to adapt to new information, it must be balanced with stability to prevent catastrophic forgetting of accumulated knowledge. To achieve this, we simulate knowledge evolution through cognitive forgetting and knowledge consolidation memory.
@@
Cognitive Forgetting Module (CFM): Improving the Expression of Historical Knowledge. In the human brain, knowledge acquisition is not a simple linear accumulation, but rather involves the active forgetting mechanism constantly clearing out outdated information to maintain the efficient operation of the cognitive system. Based on Ebbinghaus’s forgetting theory (Ebbinghaus 1985), we have introduced a cognitive forgetting module to process view information from past moments. This theory states that memory decay follows a pattern of rapid decline initially, followed by a slower rate, so the contribution of historical knowledge should not be treated equally but should gradually diminish over time. Since the mathematical properties of power-law functions exhibit a trend of rapid decline followed by gradual reduction (Heathcote, Brown, and Mewhort 2000), we utilize this tool to model the decay process. Specifically, for past representation $\mathbf{Z}_{i}$ (where $i\in\{1,...,t-1\}$), we calculate weight $w_{i}^{(t)}$ based on its temporal relevance to the current step $t$ by power-law functions:
$w_{i}^{(t)}=\frac{(t-i)^{-\lambda}}{\sum_{j=1}^{t-1}(t-j)^{-\lambda}},\quad i\in\{1,\dots,t-1\}.$ (5)
@@
CFM assigns higher weights to recent representations and allows the contributions of earlier representations to decay. This scheme simulates the pattern of memory evolution in the brain to improve the expression of historical knowledge.
Knowledge Consolidation Memory Module (KCM): Integration for Stable Long-Term Memory. To address the critical challenge of stability in continuous learning, we draw inspiration from the memory mechanism of the prefrontal cortex. In the human brain, the prefrontal cortex collaborates with the hippocampus to convert short-term knowledge into stable long-term memory. To simulate this collaborative mechanism, we propose a knowledge consolidation memory module (KCM). Specifically, we construct a temporal tensor $\mathcal{Z}\in\mathbb{R}^{n\times m\times 2}$ by stacking the historical representation $\mathbf{Z}_{hist}$ and the current representation $\mathbf{Z}_{t}$, thereby achieving high-order interaction between old and new knowledge. This process is expressed as
$\mathcal{Z}=stack(\mathbf{Z}_{hist},\mathbf{Z}_{t},3).$ (7)
where $\text{stack}(\cdot,\cdot,3)$ denotes stacking along the third dimension (where the third dimension has only two slices, among which $\mathcal{Z}(:,:,1)=\mathbf{Z}_{hist},\mathcal{Z}(:,:,2)=\mathbf{Z}_{t},$). Next, to enhance consistency and consolidate the merged knowledge into a low-rank structure, we applied alternative rank minimization regularization (ARMR), which is defined in Definition 2. The process is expressed as
@@
KCM combines historical and current representations into a higher-order tensor to simulate the integration of the prefrontal cortex. Meanwhile, low-rank approximation captures the most critical structural information in the data, aligning with the goal of memory consolidation. This process correlates with the properties of the prefrontal cortex, ultimately forming a stable representation of long-term knowledge.
The Overall Objective Function
By combining the above modules responsible for plasticity and stability, the optimization objective of MemEvo can be expressed as:
$\mathcal{L}_{MemEvo}=\mathcal{L}_{recon}+\alpha\mathcal{L}_{align}+\beta\mathcal{L}_{consolidate}.$ (9)
The final optimization problem is then formulated as

### v2

# Association and Consolidation: Evolutionary Memory-Enhanced Incremental Multi-View Clustering

@@ Abstract
Incremental multi-view clustering aims to achieve stable clustering results while addressing the stability-plasticity dilemma (SPD) in view-incremental scenarios. The core challenge is that the model must have enough plasticity to quickly adapt to new data, while maintaining sufficient stability to consolidate long-term knowledge. To address this challenge, we propose a novel Evolutionary Memory-Enhanced Incremental Multi-View Clustering (EMIMC), inspired by the memory regulation mechanisms of the human brain. Specifically, we design a rapid association module to establish connections between new and historical views, thereby ensuring the plasticity required for learning new knowledge. Second, a cognitive forgetting module with a decay mechanism is introduced. By dynamically adjusting the contribution of the historical view to optimize knowledge integration. Finally, we propose a knowledge consolidation module to progressively refine short-term knowledge into stable long-term memory using temporal tensors, thereby ensuring model stability. By integrating these modules, EMIMC achieves strong knowledge retention capabilities in scenarios with growing views. Extensive experiments demonstrate that EMIMC exhibits remarkable advantages over existing state-of-the-art methods.
1 Introduction
@@
Recent advancements in multimedia technology have significantly transformed data analysis paradigms, where Multi-View Clustering (MVC) has demonstrated notable advantages in unsupervised scenarios by integrating heterogeneous data sources [33, 8, 37, 11]. Traditional MVC methods have achieved significant success in processing static, complete datasets, as they require a complete set of views to achieve effective knowledge fusion [7]. However, data is not static but evolves dynamically over time in many real-world applications. For example, medical monitoring may extend from clinical measurements to imaging scan data, and multi-view data facilitates tracking of patient progress. This view-incremental property renders traditional MVC methods inefficient or even infeasible, as retraining the model from scratch with each new view fails to leverage previously learned knowledge. Consequently, Incremental Multi-View Clustering (IMVC) has become a critical research direction [6, 35, 10, 32], which aims to update clustering results efficiently as new views arrive. However, its core challenge is to resolve a fundamental contradiction: the Stability-Plasticity Dilemma (SPD) [38, 17]. On the one hand, the model must possess sufficient plasticity to enable rapid learning and adaptation to new patterns and knowledge contained in new data. If the model is too rigid, it will be unable to capture dynamic changes in the data, leading to a rapid decline in performance. On the other hand, the model must also maintain sufficient stability to consolidate long-term accumulated knowledge and avoid catastrophic forgetting when learning new information. Therefore, an ideal IMVC model must achieve a balance between these two aspects.
@@
Recent incremental multi-view clustering (IMVC) methods have shown significant progress. For example, Wan et al. [32] update a consensus partition matrix by transferring knowledge from historical views, enabling incremental learning. Yan et al. [39] further introduced a category memory library to store learned categories, while Qu et al. [28] proposed a lightweight framework that retains and jointly optimizes the consensus anchor graph of previous views. Despite existing IMVC methods having made valuable explorations, they often address only specific aspects of SPD. Some prioritize plasticity, rapidly integrating new data at the risk of catastrophic forgetting, while others emphasize stability through regularization, potentially hindering adaptation to new views. A key limitation is that existing methods fail to balance old knowledge retention and new knowledge integration, which causes performance bottlenecks as views continuously grow. To address this dilemma, we drew inspiration from the memory system of the human brain. Cognitive science research indicates that the collaborative memory system between the hippocampus and the prefrontal cortex serves as a paradigm for balancing plasticity and stability. The hippocampus can quickly encode new episodic memories and is highly plastic [2]. In contrast, the prefrontal cortex integrates and consolidates these temporary memories into highly stable long-term knowledge through a slow process. This complementary mechanism enables humans to continue learning without easily forgetting the past. Furthermore, knowledge acquisition is not merely information accumulation. The brain also employs a complex active adaptive forgetting mechanism to discard outdated information and maintain cognitive efficiency [1]. This forgetting is not a passive process of memory decay but an active process of neural remodeling that occurs immediately after learning behavior is completed [13].
@@
Inspired by these cognitive principles, we propose an Evolutionary Memory-Enhanced Incremental Multi-View Clustering model (EMIMC) designed to address the SPD. We should clarify that EMIMC does not aim for biologically accurate simulation. Instead, it draws functional inspiration from the brain’s effective division of labor between plastic learning and stable integration to construct an effective algorithmic. To this end, we design a collaborative framework consisting of three core modules. First, a hippocampus-inspired rapid association module establishes connections between new and historical views, ensuring the model’s plasticity for new knowledge learning. Second, we introduce a Cognitive Forgetting Module (CFM) that incorporates a decay mechanism functionally similar to human active forgetting, optimizing knowledge fusion by dynamically adjusting historical contributions. Finally, we design a prefrontal cortex-inspired Knowledge Consolidation Module (KCM), which leverages the stability of the temporal tensor to refine short-term knowledge into stable long-term memory. By integrating these three functionally complementary modules, EMIMC achieves remarkable knowledge retention capabilities in view-incremental scenarios. Our main contributions are summarized as follows:
@@
  • EMIMC is the first to introduce the functional principles of the hippocampus-prefrontal cortex collaborative memory model to incremental multi-view clustering, providing a new paradigm to address the Stability-Plasticity Dilemma.
  • We design novel cognitive forgetting and knowledge consolidation modules. The former mitigates catastrophic forgetting by dynamically adjusting historical weights, while the latter utilizes temporal tensors to mathematically instantiate the refinement of short-term knowledge into stable long-term memory.
@@
  • Extensive experiments show that our proposed method outperforms existing incremental clustering methods.
2 Preliminaries
2.1 Incremental multi-view clustering
@@
Incremental Multi-View Clustering (IMVC) continuously integrates new views while retaining historical knowledge, thereby achieving stable clustering. Existing literature achieves incremental updates by maintaining historical knowledge carriers  [32, 23, 34, 6]. For example, CMVC [32] and CAC [39] maintain consensus partition matrices to achieve knowledge reuse. FDMVC  [23] maintains semantically consistent representations for updates. LAIMVC  [28] maintains consensus anchor graphs and aligns new view anchors through permutation matrices. Consistent with existing research, EMIMC is based on the following assumptions: each view contains complete class information, and class growth scenarios are not considered. Based on this, we focus on drawing functional inspiration from cognitive mechanisms to achieve an efficient capture of multi-view information.
2.2 Tensor for knowledge consolidation
In technical implementation, a core challenge of multi-view clustering is fusing information from diverse views to obtain consensus representations. Recently, representing multi-view data as tensor and leveraging low-rank constraints to capture their underlying high-order correlations has become the mainstream technical paradigm in this field  [5, 12, 21, 15]. This paradigm provides powerful mathematical tools for modeling knowledge consolidation. However, minimizing the tensor rank is a nonconvex and NP-hard problem  [44, 27]. To address this challenge, researchers have proposed nonconvex alternatives, which can more efficiently approximate the tensor rank. Let $\{\mathbf{X}_{t}\}_{t=1}^{V}\in\mathbb{R}^{n\times d_{t}}$ be multi-view data with $V$ views, where $d_{t}$ is the dimension of the view and $n$ is the sample number. We can optimize the following objective function to achieve our goal.
$\min_{\mathbf{A}_{t},\mathbf{Z}_{t}}\sum_{t=1}^{V}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1}+\lambda\|\mathcal{Z}\|_{low},s.t.\mathcal{Z}=\Phi(\mathbf{Z}_{1},..,\mathbf{Z}_{t}).$ (1)
@@ 2.2 Tensor for knowledge consolidation
where $\mathbf{A}_{t}\in\mathbb{R}^{m\times d_{t}}$ ($m$ is the latent space dimension) is the basic matrix and $\mathbf{Z}_{t}$ is the representation for $t$-view. Besides, the function $\Phi(\cdot)$ constructs a third-order tensor $\mathcal{Z}\in\mathbb{R}^{n\times m\times V}$ by stacking the representations $\mathbf{Z}_{t}$ from each view, $\|\cdot\|_{low}$ is the rank constraint term. Here, a commonly used nonconvex substitution function  [22] is introduced, which is defined as:
@@
$\|\mathcal{Z}\|_{ARMR}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\sum_{i=1}^{min(n_{1},n_{2})}\left(\frac{1-e^{-\mathcal{S}_{f}^{k}(i,i)}}{1+e^{-\mathcal{S}_{f}^{k}(i,i)}}\right).$ (2)
@@
With the Fast Fourier Transform (FFT) along the third dimension, we can obtain $\mathcal{Z}_{f}^{k}$ by $fft(\mathcal{Z})=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\mathcal{Z}_{f}^{k}$. Then, we perform tensor singular value decomposition(t-SVD) to obtain $\mathcal{Z}_{f}^{k}=\mathcal{U}*\mathcal{S}_{f}^{k}*\mathcal{V}^{T}$[16].
3 The Proposed method
@@
Figure 1: The framework of EMIMC. The model achieves the retention of historical knowledge and the learning of new knowledge through three core modules: Rapid Associative, Cognitive Forgetting, and Knowledge Consolidation Module.
3.1 Overview and Motivation
@@
In incremental multi-view clustering, maintaining a balance between learning new knowledge (plasticity) and retaining existing knowledge (stability) is a major challenge. To address this, we draw functional inspiration from the human brain’s collaborative memory mechanism between the hippocampus and the prefrontal cortex. As shown in Figure 1, EMIMC contains three core components functionally designed to address the Stability-Plasticity Dilemma.
3.2 Plasticity: New Knowledge Acquisition
@@
When new view data arrives, the model must quickly adapt to the new knowledge. We functionally emulate the brain’s rapid learning ability through incremental representation learning and a hippocampus-inspired rapid associative mechanism.
@@
1) Foundation: Incremental Representation Learning. First, we design an incremental representation learning strategy. When a new data view $\mathbf{X}_{t}$ is received, we reconstruct it using a representation $\mathbf{Z}_{t}$ and an orthogonal basis matrix $\mathbf{A}_{t}$:
$\mathcal{L}_{recon}=\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1},\quad s.t.\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I},$ (3)
@@
where $\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I}$ ensures uncorrelated basis vectors [4], and the $\ell_{2,1}$ norm enhances robustness [26].
@@
2) Rapid Associative Module (RAM). In cognitive science, a key function of the hippocampus is associative memory, which rapidly establishes connections between newly received information and recent information [25]. To simulate this rapid associative capability, we propose a hippocampus-inspired Rapid Associative Module (RAM). This module aims to facilitate immediate interaction between new and old knowledge, enabling the model to adapt to historical contexts while learning new tasks. Specifically, we introduce an orthogonal matrix $\mathbf{P}_{t}\in\mathbb{R}^{m\times m}$ to connect the current representation $\mathbf{Z}_{t}$ with the previous representation $\mathbf{Z}_{t-1}$. This process can be expressed as:
$\mathcal{L}_{associate}=\|\mathbf{Z}_{t}-{\mathbf{Z}_{t-1}\mathbf{P}_{t}}\|_{F}^{2},s.t.\mathbf{P}_{t}\mathbf{P}_{t}^{T}=\mathbf{I}.$ (4)
Although RAM employs a technically mature orthogonal method, this choice is a principled abstraction of the hippocampus. This function simulates the rapid associative processes occurring in the hippocampus: new memories must establish connections with existing memories during the learning process, thereby enabling rapid acquisition of new knowledge.
3.3 Stability: Knowledge Consolidation and Evolution
@@
Although plasticity enables systems to adapt to new information, it must be balanced with stability to prevent catastrophic forgetting of accumulated knowledge. To achieve this, we functionally simulate this process through cognitive forgetting and the knowledge consolidation memory module.
1) Cognitive Forgetting Module (CFM).
@@
In the human brain, knowledge acquisition is not a simple linear accumulation, but rather involves an active forgetting mechanism. This mechanism continuously removes outdated information to maintain the efficient functioning of the cognitive system. Based on Ebbinghaus’s forgetting theory [9], we have introduced a cognitive forgetting module to process view information from past moments. According to this theory, memory decay is initially rapid and then slows down. Therefore, historical knowledge should not contribute equally; instead, its influence should progressively diminish as new views are introduced. Since the mathematical properties of power-law functions exhibit a trend of rapid decline followed by gradual reduction [14], we utilize this tool to model the decay process. Specifically, for past representation $\mathbf{Z}_{i}$ (where $i\in\{1,2,...,t-1\}$), we compute the weight $w_{i}^{(t)}$ using a power-law function based on its relevance to the current view $t$:
$w_{i}^{(t)}=\frac{(t-i)^{-\lambda}}{\sum_{j=1}^{t-1}(t-j)^{-\lambda}},\quad i\in\{1,2,\dots,t-1\},$ (5)
@@
CFM assigns higher weights to recent representations and decays the contributions of earlier representations. This mechanism functionally mimics the active memory patterns, thereby enhancing historical knowledge expression.
2) Knowledge Consolidation Memory Module (KCM).
To address long-term stability issues, we draw inspiration from the prefrontal cortex’s mechanism of integrating short-term memories into stable patterns. Standard tensor methods typically stack all views into an ever-growing tensor, leading to a cumulative increase in computational complexity. We propose the Knowledge Consolidation Memory (KCM) module. In contrast, our module draws inspiration from the prefrontal cortex’s mechanism of integrating interactions between long-term memory and current memory. Specifically, by superimposing the historical representation $\mathbf{Z}_{hist}$ with the current representation $\mathbf{Z}_{t}$, we construct a temporal tensor $\mathcal{Z}\in\mathbb{R}^{n\times m\times 2}$ to achieve higher-order interaction between old and new knowledge:
$\mathcal{Z}=stack(\mathbf{Z}_{hist},\mathbf{Z}_{t},3),$ (7)
where $\text{stack}(\cdot,\cdot,3)$ denotes stacking along the third dimension (where the third dimension has only two slices, among which $\mathcal{Z}(:,:,1)=\mathbf{Z}_{hist},\mathcal{Z}(:,:,2)=\mathbf{Z}_{t}$). Next, we need to simulate how the prefrontal cortex consolidates this information into long-term knowledge. To achieve this computationally, we employ the low-rank tensor constraint paradigm discussed in Definition 2. We treat the tensor $\mathcal{Z}$, formed by stacking historical and current representations, as a high-order knowledge repository. By imposing low-rank constraints on it, we force the model to retain only the most essential structural information across time, functionally simulating the process of knowledge consolidation:
@@
3.4 The Overall Objective Function
By combining the functionally inspired modules responsible for plasticity and stability, the overall objective function of EMIMC is formulated as:
$\mathcal{L}_{EMIMC}=\mathcal{L}_{recon}+\alpha\mathcal{L}_{associate}+\beta\mathcal{L}_{consolidate}.$ (9)
The final optimization problem is formally stated as: